# An Efficient Finite Difference Approximation via a Double Sample-Recycling Approach

### v1

@@ 6 Numerical Experiments
  • Bias: The bias is influenced by the signs of $B$ and $D$. If $B$ and $D$ have the same (opposite) signs, the absolute bias of the DSR-CFD method is smaller (larger) than that of the Opt-CFD method. This aligns with Theorem 4 because $\bm{c^{\top}Pc^{4}}$ is negative. When $B$ and $D$ share the same (opposite) signs, $B$ and $D\bm{c^{\top}Pc^{4}}$ exhibit opposite (same) signs. From Proposition 5, the ratio of the abstract values of biases of the DSR-CFD estimator and the Opt-CFD estimator is bounded by 1 or $1+\sqrt{\frac{r}{K}}\frac{2}{\sigma(\theta_{0})}\left|D{\bm{c^{\top}Pc^{4}}}\right|$.
[…]
  • Variance: It is observed that as the sample size increases, the variance of the DSR-CFD estimator consistently outperforms that of the Opt-CFD estimator, aligning with Proposition 5. For instance, at $x=1$ with a sample budget of $10^{5}$, the variances for the DSR-CFD and Opt-CFD estimators are $9.8655\times 10^{-4}$ and $0.0015$, respectively.
  • MSE: It is important to note that the bias and variance of the DSR-CFD estimator consistently approximate and are smaller than those of the Opt-CFD estimator, respectively. Consequently, when employing MSE as the error criterion, the DSR-CFD estimator consistently performs nearly as well as (or even outperforms) the Opt-CFD estimator.
We examine the estimated MSE in Example 2 by setting $\theta_{0}=0$ and varying the sample pairs from $n=50$ to $n=500$. When considering the EM-CFD algorithm, we set $r=0.1$ and generate $h_{1},...,h_{rn}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,0.01\times(rn)^{-1/5}\right)$. When applying the DSR-CFD algorithm, we set $K=10$ and $r=1$. That is, the entire simulation budget is allocated to the pilot phase. Figure 4 shows the results. From the figure, we observe that the estimated MSE of the EM-CFD estimator decreases gradually as the sample size increases. Compared to the Opt-CFD and DSR-CFD estimators, however, the EM-CFD estimator exhibits a significantly larger MSE. Moreover, the MSE of the DSR-CFD estimator is consistently close to that of the Opt-CFD estimator. For example, when the sample size is 50, the MSE for the EM-CFD method is approximately 5.959, while the MSEs for the DSR-CFD and Opt-CFD methods are approximately 0.421 and 0.170, respectively.
Example 3.
We consider a generic $M/M/1$ queueing system which is empty initially. Both of the arrival distribution and the service distribution are the exponential distribution and the arrival rate and service rate are denoted by $\lambda$ and $\mu$, respectively. Observations are the averaged system time of the first $N$ customers and we are interested in the gradient of this quantity w.r.t. the service rate. For this example, we consider two cases.
  • Case 1: We consider a critically loaded system, where we set $\lambda=\mu=4$ and $N=10$. In this case, the true derivative is $-0.2501$, which is calculated using the likelihood ratio/score function method (see e.g. Glynn, 1990) with $10^{6}$ simulation repetitions (Lam et al., 2022).
@@
In this example, we denote the estimate of $B$ as $\widetilde{B}$ in Appendix B.1 and set $\epsilon=5\times 10^{-4}$. Figure 5 illustrates the numerical results for different methods in Example 3 when the sample-pair size changes from $n=60$ to $n=1000$. We compare the following three methods: The conventional central finite difference (Con-CFD) method, EM-CFD method and DSR-CFD method. For the Con-CFD method, we do not know any information about the model in advance and set $B=5$ and $\sigma^{2}(\theta_{0})=1$. When considering the EM-CFD method, we set $r=0.1$ and generate $h_{1},...,h_{rn}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,1\times(rn)^{-1/5}\right)$. When applying the DSR-CFD method, we set $K=20$ in the pilot phase and $r=1$.
@@
Figure 5 illustrates that the performance of the Con-CFD estimator is inferior compared to the other two estimators. This observation emphasizes the importance of estimating unknown constants within finite difference methods. In case 1 of Example 3, the DSR-CFD estimator consistently yields the most favorable results. For instance, with a sample-pair size of 60, the MSEs for the DSR-CFD and EM-CFD estimators are $0.011$ and $0.033$, respectively. In case 2, the performances of these two estimators are similar. In conclusion, the DSR-CFD estimator performs effectively across various problems.
@@ 7 Conclusions
The choice of the perturbation significantly impacts the accuracy of the finite difference approximation in stochastic gradient estimation. Although the order of the optimal perturbation is known, determining the unknown constant remains a challenge. To address this issue and construct an efficient finite difference approximation, in this paper, we have developed a double sample-recycling approach aimed at maximizing sample efficiency. Theoretical analyses reveal that the proposed estimator exhibits a reduced variance compared to the optimal finite difference estimator, and in some cases, a decrease in bias. In particular, we provide algorithms and numerical experiments. The theory and the findings of the numerical experiments are consistent.
@@ A.3 Asymptotic property of $\widehat{\alpha}^{\prime}(\theta_{0})$
The proof of Corollary 6 is identical to that of Theorem 2. (59) provides the convergence rates for the bias and the variance of $\widehat{\alpha}^{\prime}(\theta_{0})$, which are $O\left(n_{b}^{4\gamma}\right)$ and $O\left(n_{b}^{-1-2\gamma}\right)$, respectively. Additionally, when $c_{1},...,c_{K}$ are generated from a proper distribution $\mathcal{P}_{0}$, the convergence rates of the bias and the variance of $\widehat{\alpha}^{\prime}(\theta_{0})$ are $O\left(n_{b}^{4\gamma}\right)$ and $O\left(n_{b}^{8\gamma}\right)+O\left(n_{b}^{-1-2\gamma}\right)$, respectively. Theorem 2 and Corollary 6 demonstrate that when estimating $B$, we also obtain a roughly estimate of $\alpha^{\prime}(\theta_{0})$, which is denoted by $\widehat{\alpha}^{\prime}(\theta_{0})$. However, directly using $\widehat{\alpha}^{\prime}(\theta_{0})$ is not advisable since it bases on a small sample size $Kn_{b}$. Nevertheless, it is evident that $\widehat{\alpha}^{\prime}(\theta_{0})$ contains valuable gradient information. To prevent the loss of the information, in Section 4, we will propose a sample-recycling-based estimator which contains $\widehat{\alpha}^{\prime}(\theta_{0})$.
@@ A.4 Analysis of (21)
$=\frac{\left(\sigma(\theta_{0})+\sigma^{\prime}(\theta_{0})h+o(h)\right)^{2}+\left(\sigma(\theta_{0})-\sigma^{\prime}(\theta_{0})h+o(h)\right)^{2}}{4h^{2}}$
$=\frac{\sigma^{2}(\theta_{0})+\sigma^{\prime}(\theta_{0})^{2}h^{2}+o(h^{2})}{2h^{2}},$
@@ B.1 Further explanations on (27)
Note that the continuous mapping theorem implies that $\left(\frac{\widehat{\bm{\beta}}_{v}}{\widehat{B}^{2}}\right)^{1/3}$ and $\left(\frac{\widehat{B}^{2}}{\widehat{\bm{\beta}}_{v}}\right)^{1/3}$ converge to $\left(\frac{\bm{\beta}_{v}}{B^{2}}\right)^{1/3}$ and $\left(\frac{B^{2}}{\bm{\beta}_{v}}\right)^{1/3}$ in probability, respectively. Note that the expectation and limit operators may not be exchangeable because 0 exists at the support set of the regression estimators. However, we don’t have to be confused about this issue. Take $\widehat{B}$ for example, assume that $B$ is bounded below away from 0. Specifically, we assume that there exists a positive number $\epsilon$ such that $B>\epsilon>0$. We can solve (14) under the constraint $\widehat{B}>\epsilon$. In others words, we denote a new estimator $\widetilde{B}=\epsilon+\max{\left\{\widehat{B}-\epsilon,0\right\}}$ and replace $\widehat{B}$ with $\widetilde{B}$. Lemma 7 and Proposition 8 indicate that the asymptotic property of $\widetilde{B}$ is identical to that of $\widehat{B}$. As a result, the dominated convergence theorem can be used to obtain the leading terms of (27).
@@
Proof: We prove (65) and (66), respectively.
@@
Proof: For a given $\delta>0$, we construct a smooth approximation function of $f(x)=\epsilon+\max{\{x-\epsilon,0\}}$, denoted by $f_{\delta}(x)$, as follows:
@@
$=$ $\underset{\delta\to 0}{\lim}f_{\delta}(B)+\underset{\delta\to 0}{\lim}f_{\delta}^{\prime}(B)H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right)$
$=$ $\epsilon+(B-\epsilon)\mathbbm{1}_{\{B>\epsilon\}}+\mathbbm{1}_{\{B>\epsilon\}}H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right)$
@@
$=$ $\underset{\delta\to 0}{\lim}f_{\delta}^{\prime}(B)^{2}V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right)$
$=$ $\mathbbm{1}_{\{B>\epsilon\}}V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right)$

### v4

@@ 6 Numerical Experiments
  • Bias: The bias is influenced by the signs of $B$ and $D$. If $B$ and $D$ have the same (opposite) signs, the absolute bias of the Cor-CFD method is smaller (larger) than that of the Opt-CFD method. This aligns with Theorem 4 because $\Lambda$ is negative.
  • Variance: It is observed that as the sample size increases, the variance of the Cor-CFD estimator consistently outperforms that of the Opt-CFD estimator, aligning with Proposition 5. For instance, at $x=3$ with a sample budget of $10^{4}$, the variances for the Cor-CFD and Opt-CFD estimators are $0.0021$ and $0.0061$, respectively.
  • MSE: It is important to note that the bias and variance of the Cor-CFD estimator consistently approximate and are smaller than those of the Opt-CFD estimator, respectively. Consequently, when employing MSE as the error criterion, the Cor-CFD estimator consistently performs nearly as well as (or even outperforms) the Opt-CFD estimator.
Example 3 (Queuing system).
In this example, we consider two cases.
  • Case 1: We consider a critically loaded system, where we set $\lambda=\mu=4$ and $N=10$. In this case, the true derivative is $-0.2501$, which is calculated using the likelihood ratio/score function method (see, e.g., Glynn, 1990) with $10^{6}$ simulation repetitions (Lam et al., 2023).
@@
Figure 3 illustrates the numerical results for different methods in Example 3 when the sample-pair size changes from $n=60$ to $n=1000$. We compare the following three methods: the traditional CFD (Tra-CFD) method, EM-CFD method and Cor-CFD method. For the Tra-CFD method, we do not know any information about the model in advance and set $B=5$ and $\sigma^{2}(\theta_{0})=1$. When considering the EM-CFD method, we set $r=0.1$ and generate $h_{1},...,h_{rn}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,1\times(rn)^{-1/5}\right)$. When applying the Cor-CFD method, we set $K=20$ and $r=1$ ($r=1$ means the total samples are pilot samples).
@@
Figure 3 illustrates that the performance of the Tra-CFD estimator is inferior to the other two estimators, underscoring the importance of estimating unknown constants in finite difference methods. In case 1 of Example 3, the Cor-CFD estimator consistently yields the most favorable results. For instance, with a sample-pair size of 60, the MSEs for the Cor-CFD and EM-CFD estimators are $0.011$ and $0.033$, respectively. In case 2, the performances of these two estimators are comparable.
7 Derivative-Free Optimization Using Cor-CFD
In this section, we consider the DFO problem, i.e., a stochastic optimization problem where gradient information is unavailable and only the stochastic function values can be obtained. Our goal is to incorporate the Cor-CFD estimator into the well-known L-BFGS algorithm (Nocedal, 1980) and numerically verify its efficiency in solving the DFO problem. We leave the corresponding theoretical studies for future research.
Consider an optimization problem
$\underset{\theta\in\mathbb{R}^{d}}{\min}\ \alpha(\theta)=\mathbb{E}[Y(\theta)].$ (45)
Suppose that Assumptions 1 and 3 hold, ensuring the applicability of our Cor-CFD method. Similarly to the operations in Shi et al. (2023), we replace the gradient estimator in the L-BFGS algorithm with our proposed Cor-CFD estimator, constructing a finite difference L-BFGS series:
$\theta_{k+1}=\theta_{k}-a_{k}H_{k}\widehat{\Delta}_{T_{k},\widehat{h}_{T_{k}}},$
where $k$ represents the iterative process, $a_{k}>0$ is the step length, $H_{k}$ is an approximation to the inverse Hessian matrix and $\widehat{\Delta}_{T_{k},\widehat{h}_{T_{k}}}$ is a Cor-CFD estimator based on $T_{k}$ sample pairs.
To ensure the seamless progress of iterations, two issues must be considered: how to determine the appropriate sample size $T_{k}$ for each iteration, and how to choose the step length $a_{k}$. For the first issue, as iterations proceed and $\theta_{k}$ approaches the optimal point, the accuracy of the finite difference estimator becomes increasingly important. Intuitively, $T_{k}$ should therefore increase as $k$ increases. The theoretical determination of the optimal $T_{k}$ could be left as a topic for future research.
For the second issue, we apply the stochastic Armijo condition (see, e.g., Berahas et al., 2019; Shi et al., 2023),
$Y\left(\theta_{k}-a_{k}H_{k}\widehat{\Delta}_{T_{k},\widehat{h}_{T_{k}}}\right)\leq Y(\theta_{k})+l_{1}a_{k}\widehat{\Delta}_{T_{k},\widehat{h}_{T_{k}}}^{\top}H_{k}\widehat{\Delta}_{T_{k},\widehat{h}_{T_{k}}}+2\sigma.$ (46)
Specifically, we shrink $a_{k}\leftarrow l_{2}a_{k}$ until the above condition holds, where $10<l_{1}<l_{2}<$ 1. Note that $\sigma$ need not to be chosen too precisely and it suffices to select it as the upper bound of $\sigma(\theta_{k})$ (Berahas et al., 2019). For example, when $\sigma$ is relatively large, the stochastic Armijo condition is naturally satisfied and $a_{k}$ degenerates a constant step length.
The Cor-CFD-based L-BFGS (CorCFD-L-BFGS) algorithm is shown in Algorithm 2. This algorithm is a quasi-Newton algorithm for DFO problems. We will compare our proposed Algorithm 2 with the traditional CFD-based L-BFGS (TraFD-L-BFGS) algorithm (Shi et al., 2023) and the new unconstrained optimization algorithm (NEWUOA) (Powell, 2006). The TraCFD-L-BFGS algorithm replaces the gradient estimator in L-BFGS algorithm by the finite difference estimator constructed from only a sample pair. Note that the NEWUOA is used as a benchmark in Shi et al. (2023). The detailed settings for each algorithm are provided Appendix D.1.
Input: The number of total sample pairs $T$, the number of perturbation parameters $K$, the number of initial batch size $T_{0}$, the number of function evluations $t=0$, the number of bootstrap $I$, initial perturbation generator $\mathcal{P}_{0}$, the line search parameters $(l_{1},l_{2})$, starting point $\theta_{0}$ and initial step length for line search $a_{0}>0$.
Initialization: Set $k=0,H_{0}=I_{d\times d}$, where $I_{d\times d}$ is an identity matrix, and $d$ is the problem dimension. Obtain the estimator $\widehat{\Delta}_{T_{0},\widehat{h}_{T_{0}}}$ using Algorithm 1.
While $t<2T$ do
  1. 1.
    Use stochastic Armijo line search (46) to find an appropriate step length $0<a_{k}<a_{0}$ and set $t=t+t_{ls}$, where $t_{ls}$ denotes the function evaluation counter during the line search procedure.
  2. 2.
    Update $\theta_{k+1}=\theta_{k}-a_{k}H_{k}\widehat{\Delta}_{T_{k},\widehat{h}_{T_{k}}}$.
  3. 3.
    Set $T_{k+1}=\lfloor(T_{k}+k+1)/K\rfloor\times K$ and obtain the estimator $\widehat{\Delta}_{T_{k+1},\widehat{h}_{T_{k+1}}}$ using Algorithm 1.
  4. 4.
    Update $H_{k+1}$ via the L-BFGS algorithm.
  5. 5.
    Set $t=t+2dT_{k}$ and $k=k+1$.
Output: The ultimate estimate $\theta_{k+1}$.
Algorithm 2 (Cor-CFD-based L-BFGS algorithm.).
We consider two stochastic optimization problems constructed by artificially imposing a standard normal noise $\mathcal{N}(0,1)$ on two classical test functions in optimization area – Rosenbrock function and Zakharov function (see,e.g., Schittkowski, 2012; Jamil and Yang, 2013) and examine the efficiency of Algorithm 2. The definitions of these two functions are shown as follows:
  • The Rosenbrock function $f:\bm{x}\mapsto f(\bm{x})$ where $\bm{x}\in\mathbb{R}^{2}$ is defined as
    $f(\bm{x})=100(x_{2}-x_{1}^{2})^{2}+(x_{1}-1)^{2}.$
  • A $d$-dimensional Zakharov function $f:\bm{x}\mapsto f(\bm{x})$ where $\bm{x}\in\mathbb{R}^{d}$ is defined as
    $f(\bm{x})=\sum_{i=1}^{d}x_{i}^{2}+\left(\sum_{i=1}^{d}0.5ix_{i}\right)^{2}+\left(\sum_{i=1}^{d}0.5ix_{i}\right)^{4}.$
That is, in (45), $Y(\theta)=f(\theta)+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,1)$.
Figure 4: Performances of the optimality gaps by different algorithms for Rosenbrock function w.r.t. different sample sizes.
We show the optimality gap (OG, i,e., $\alpha(\theta_{k})-\alpha(\theta^{*})$, where $\theta^{*}$ is the global minimum point) for different algorithms on the Rosenbrock function in Figure 4. Figure 4 reveals that the TraCFD-L-BFGS algorithm fails to converge due to the large error in gradient estimation when using only one sample pair, which prevents the algorithm from identifying the correct descent direction. In contrast, the CorCFD-L-BFGS algorithm employs a batch of samples to estimate the gradient, facilitating the algorithm’s gradual convergence to the minimum point. In addition, compared to NEWUOA, the CorCFD-L-BFGS algorithm also produces better outputs. For example, when the sample-pair size is $10^{5}\times d$, the optimal gaps for the CorCFD-L-BFGS algorithm and NEWUOA are 0.12 and 0.61, respectively.
Note that the TraCFD-L-BFGS algorithm is slightly better when the sample size is small. This may be attributed to two factors. Firstly, the TraCFD-L-BFGS algorithm undergoes more iterations than the CorCFD-L-BFGS algorithm because the finite difference estimator in the former is constructed by only one sample pair. Secondly, in the initial a few iterations, the gradient descent through the rough gradient estimation may be more important than the accuracy of the gradient descent direction. However, when the sample-pair size is relatively larger, i.e., $2\times 10^{3}$, the CorCFD-L-BFGS algorithm outperforms the TraCFD-L-BFGS algorithm, and so the former is more recommended.
Next, we study the comparison among these algorithms on the Zakharov function. Table 4 exhibits the solution gap (SG, denoted by $||\theta_{k}-\theta^{*}||_{2}$) and OG related to these algorithms. It is worth mentioning that in this example, the starting point $\theta_{0}=(\underbrace{1,...,1}_{d})^{\top}$, and the corresponding function values are about $f(\theta_{0})=1.3125$, $5.7191\times 10^{5}$ and $4.0649\times 10^{13}$ for $d=1,10$ and 100, respectively. The optimal point and the optimal function value are $\theta^{*}=0$ and $f(\theta^{*})=0$, respectively. In the table, “$\approx 0$” indicates a value less than $10^{-3}$.
Table 4: Performances of the solution gaps and optimality gaps by different algorithms for Zakharov function w.r.t. different dimensions and different sample sizes.
Dimension $d=1$ $d=10$ $d=100$
Sample Size (×$\times d$) $10^{3}$ $10^{4}$ $10^{5}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{3}$ $10^{4}$ $10^{5}$
NEWUOA (SG) 0.059 $\approx 0$ $\approx 0$ 5.654 7.542 7.532 8.254 32.03 54.09
TraCFD-L-BFGS (SG) 0.087 0.081 0.079 3.591 3.684 3.697 7.718 14.92 17.20
CorCFD-L-BFGS (SG) 0.002 0.002 $\approx 0$ 0.330 0.099 0.056 4.930 1.825 0.694
NEWUOA (OG) 0.018 $\approx 0$ $\approx 0$ 30.93 57.38 57.25 68.51 1027 2927
TraCFD-L-BFGS (OG) 0.015 0.013 0.012 37.33 34.74 36.23 688.4 1469 1015
CorCFD-L-BFGS (OG) 0.004 0.002 $\approx 0$ 0.176 0.037 0.022 24.42 3.527 0.569
From Table 4, it can be seen that in the case of low dimension, e.g., $d=1$, the results of three algorithms are comparable in the sense of both SG and OG. All of them are close to 0. When the dimension is higher, e.g., $d=10$ and 100, the results of our proposed CorCFD-L-BFGS algorithm exhibit convergence as the sample-pair size increases, while those of the other algorithms do not. This is because the TraCFD-L-BFGS algorithm cannot find the correct gradient descent direction due to rough gradient estimation with only one sample pair, and the performance of NEWUOA may be improved through finding better algorithm parameter values tailored to the higher dimension cases. The detailed discussion on this topic will be provided in Appendix D.2. In particular, when $d=100$ and the sample-pair size is $10^{3}\times d$, our proposed CorCFD-L-BFGS algorithm reduces the estimated optimal function value to 24.42, down from the initial value of $4.0649\times 10^{13}$. Furthermore, with an increased sample-pair size of $10^{5}\times d=10^{7}$, the estimated optimal function value decreases to 0.569. It is worth mentioning that this sample-pair size, i.e., $10^{7}$ is considered acceptable for DFO problems with dimensions as large as $d=100$ (see, e.g., Bollapragada et al., 2024).
In addition, KW and SPSA algorithms are classical approaches in simulation optimization, a direct comparison with our proposed CorCFD-L-BFGS algorithm may not be entirely equitable. This is because KW and SPSA correspond to the first-order method, whereas CorCFD-L-BFGS corresponds to the second-order method. Nevertheless, for the sake of completeness, we provide a brief comparison and discussion of these methods in Appendix D.3. A more comprehensive and in-depth comparative analysis of these algorithms could be left as a topic for future research.
@@ 8 Conclusions
The choice of the perturbation significantly influences the accuracy of the FD estimator in stochastic gradient estimation. Although the order of the optimal perturbation is known, determining the unknown constants remains challenging. To address this issue, we have proposed a sample-driven method. Furthermore, we reuse samples, thereby inducing correlated samples to construct an efficient FD estimator. This estimator has an asymptotic variance reduction and possible bias reduction compared to the traditional optimal FD estimator. Numerical results demonstrate that our proposed correlation-induced FD estimator performs well even with small sample sizes. In addition, we propose an algorithm that combines our estimator with the renowned L-BFGS algorithm to solve DFO problems. Numerical results show that the proposed algorithm performs effectively when solving DFO problems with dimensions as large as 100.
For future work, bias reduction and variance reduction techniques for our proposed finite difference estimator could be explored. Section 7.1.2 in (Glasserman, 2013) reviews an extrapolation method for reducing the bias of conventional finite difference estimators, which could potentially be applied to our proposed estimator to achieve similar bias reduction.
On the other hand, as systems become increasingly complex, it is often easier to experiment with them rather than to fully understand their intricacies (Golovin et al., 2017). Consequently, DFO is gaining importance, and FD approximation emerges as a powerful technique for solving DFO problems (Shi et al., 2023). However, this research area remains underdeveloped. In this paper, we establish an efficient FD estimator for practical implementation and propose Algorithm 2 to solve DFO problems. In this research direction, the first topic is to prove the convergence and convergence rate of this algorithm, which intuitively, can be derived from Theorem 1 in Hu and Fu (2024). The second one involves determining the optimal sample size $T_{k}$ in each iteration (e.g., $T_{k}$ in the step 3 of Algorithm 2). The third one is how to apply the proposed FD estimator to DFO problems with large dimensions, such as $10^{4}$ or larger. A potential approach could be the combination of the FD estimator with coordinate descent methods (see e.g., Nesterov, 2012).
@@ A.3 Asymptotic Property of $\widehat{\alpha}^{\prime}(\theta_{0})$
The proof of Corollary 6 is identical to that of Theorem 2. (63) provides the convergence rates for the bias and the variance of $\widehat{\alpha}^{\prime}(\theta_{0})$, which are $O\left(n_{b}^{4\gamma}\right)$ and $O\left(n_{b}^{-1-2\gamma}\right)$, respectively. Additionally, when $c_{1},...,c_{K}$ are generated from a proper distribution $\mathcal{P}_{0}$, the convergence rates of the bias and the variance of $\widehat{\alpha}^{\prime}(\theta_{0})$ are $O\left(n_{b}^{4\gamma}\right)$ and $O\left(n_{b}^{8\gamma}\right)+O\left(n_{b}^{-1-2\gamma}\right)$, respectively.
@@ A.4 Analysis of (21)
$=$ $\frac{\left(\sigma(\theta_{0})+\sigma^{\prime}(\theta_{0})h+o(h)\right)^{2}+\left(\sigma(\theta_{0})-\sigma^{\prime}(\theta_{0})h+o(h)\right)^{2}}{4h^{2}}=\frac{\sigma^{2}(\theta_{0})+\sigma^{\prime}(\theta_{0})^{2}h^{2}+o(h^{2})}{2h^{2}},$
@@ B.1 Further Explanations on (27)
We need to demonstrate that the leading terms and $o_{p}(\cdot)$ terms in (26) converge to their corresponding terms in (27). To prevent excessive technical complexity of proving the convergence of the $o_{p}(\cdot)$ terms in (26), we adopt the operation in Hong et al. (2017) and thus directly assume that their uniform integrability holds, which ensures their convergence. Next, we prove the convergence of the leading terms in (26).
The continuous mapping theorem implies that $\left({\widehat{\bm{\beta}}_{v}}/{\widehat{B}^{2}}\right)^{1/3}$ and $\left({\widehat{B}^{2}}/{\widehat{\bm{\beta}}_{v}}\right)^{1/3}$ converge to $\left({\bm{\beta}_{v}}/{B^{2}}\right)^{1/3}$ and $\left({B^{2}}/{\bm{\beta}_{v}}\right)^{1/3}$ in probability, respectively. To prove the convergence of the leading terms in (26), we aim to apply the dominated convergence theorem. This theorem requires that $\left({\widehat{\bm{\beta}}_{v}}/{\widehat{B}^{2}}\right)^{1/3}$ and $\left({\widehat{B}^{2}}/{\widehat{\bm{\beta}}_{v}}\right)^{1/3}$ are bounded by some integrable random variable. This requirement is not apparently satisfied because 0 cannot be excluded from the support sets of the regression estimators $\widehat{\bm{\beta}}_{v}$ and $\widehat{B}$. In the following, we provide a technique, inspired by Hong et al. (2017), to prove the convergence of the leading terms in (26).
Take the convergence of the bias term in (26) for example. Note that $B\neq 0$. Specifically, we can assume that there exists a positive number $\epsilon$ such that $B>\epsilon>0$. Then, we can solve (14) under the constraint $\widehat{B}>\epsilon$. In others words, define a new estimator $\widetilde{B}=\epsilon+\max{\left\{\widehat{B}-\epsilon,0\right\}}$ and replace $\widehat{B}$ with $\widetilde{B}$ in the bias term. The following Lemma 1 and Proposition 7 demonstrate that the asymptotic properties of $\widetilde{B}$ are identical to those of $\widehat{B}$. As a result, the dominated convergence theorem ensures the convergence of the leading terms in (26) to their corresponding terms in (27).
@@
Proof.
We prove (69) and (70), respectively.
@@
Proof.
For a given $\delta>0$, we construct a smooth approximation function of $f(x)=\epsilon+\max{\{x-\epsilon,0\}}$, denoted by $f_{\delta}(x)$, as follows:
@@
$=$ $\underset{\delta\to 0}{\lim}f_{\delta}(B)+\underset{\delta\to 0}{\lim}f_{\delta}^{\prime}(B)H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right)=\epsilon+(B-\epsilon)\mathbbm{1}_{\{B>\epsilon\}}+\mathbbm{1}_{\{B>\epsilon\}}H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right)$
@@
$=$ $\underset{\delta\to 0}{\lim}f_{\delta}^{\prime}(B)^{2}V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right)=\mathbbm{1}_{\{B>\epsilon\}}V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right)$